# Optimizing GPT for Video Understanding: Zero-Shot Performance and Prompt Engineering

### v2

@@ Abstract
In this study, we tackle industry challenges in video content classification by exploring and optimizing GPT-based models for zero-shot classification across seven critical categories of video quality. We contribute a novel approach to improving GPT’s performance through prompt optimization and policy refinement, demonstrating that simplifying complex policies significantly reduces false negatives. Additionally, we introduce a new decomposition-aggregation-based prompt engineering technique, which outperforms traditional single-prompt methods. These experiments, conducted on real industry problems, show that thoughtful prompt design can substantially enhance GPT’s performance without additional finetuning, offering an effective and scalable solution for improving video classification systems across various domains in industry.
@@ 1 Introduction
In the past decade, recognition over multi-modalities has become an increasingly critical challenge for large video platform companies, such as TikTok and YouTube, which rely on managing vast amounts of user-generated content. Unlike traditional classification tasks that process single-modality inputs, multimodal classification combines multiple data sources, such as image, audio, and text, to provide more accurate and context-aware predictions. The ability to reason across these different modalities is essential for effective content classification. This capability is particularly important for identifying inappropriate content, or videos that do not align with platform policies, as it ensures a more nuanced understanding of content beyond just visual or textual cues.
@@
With the advent of large language models (LLMs) like GPT-4, a new paradigm has emerged in the field of multimodal classification. LLMs offer the promise of generalization, boasting the ability to perform a wide range of tasks with minimal task-specific data, often through one-shot or few-shot learning. Their large-scale pretraining on diverse datasets imbues them with a form of “world knowledge” that can be flexibly adapted to a range of problems, including those that involve multimodal inputs. As a result, generative AI models are now being explored as a promising solution for problems traditionally addressed by multimodal architectures. One domain ripe for this exploration is video classification, where the challenge lies in understanding video content through a combination of visual, auditory, and sometimes textual cues.
@@
However, while LLMs like GPT-4 offer broad generalization, applying them to real-world, industry-specific problems has proven difficult. These models are typically pretrained on massive public datasets, meaning that company-specific data—especially multimodal data related to proprietary video content—is rarely seen during pretraining. As such, LLMs may struggle to capture the nuances of these specific tasks without further tuning. Additionally, despite the hype surrounding generative models, there is still skepticism in the industry regarding their return on investment.
@@
In this study, we aim to address these challenges by evaluating the performance of GPT-4 in real-world, industry-relevant video classification tasks that require multimodal understanding. Specifically, we explore the potential of GPT-4o for classifying TikTok video content based on a variety of criteria. Our work differs from traditional multimodal models in two key ways: first, by focusing on the zero-shot and few-shot capabilities of GPT-4, we investigate whether large, pretrained generative models can be effectively applied to multimodal tasks without extensive fine-tuning. Second, we explore the impact of policy refinement and prompt engineering on GPT-4’s performance, addressing one of the core limitations of LLMs in industry—namely, their difficulty in handling complex, highly specific policies like those used to moderate video content.
@@
Our contributions are twofold. First, we present a comprehensive evaluation of GPT-4’s ability to handle real-world, multimodal video classification tasks, comparing its performance to existing, in-house classification models that leverage multimodal features. We show that GPT-4 can perform on par with these models in several categories, but faces challenges in more complex classifications, such as clickbait videos. Second, we introduce novel techniques to enhance GPT-4’s performance in these tasks. We find that shortening complex policy prompts improves GPT-4’s ability to classify videos accurately by reducing false negatives, and that prompt engineering—specifically, dividing tasks like clickbait detection into subcategories—can lead to significant performance gains. Another highlight is our experiments are grounded in actual industry datasets and problems, making the results highly applicable to video classification tasks at scale.
@@ 2 Related Work
Multi-modal Classification. Historically, multimodal modeling has been approached using modality-specific architectures. For instance, BERT (Devlin, 2018), RoBERTa (Liu, 2019), ResNet (He et al., 2016), and ViT (Dosovitskiy, 2020) are commonly employed for text and vision tasks, respectively. Multi-modal models like CLIP (Radford et al., 2021) and GLIP (Li et al., 2022) have effectively combined visual and textual inputs, leveraging correlations between modalities to achieve strong classification and detection results. However, these models typically require extensive task-specific fine-tuning and are limited by the rigid nature of their underlying architectures. This has resulted in a fragmented landscape where models are highly specialized for narrow tasks and lack generalizability to unseen problems or broader domains without significant retraining.
@@ 3.1 Problem Setup
Two major challenges to this problem are: (1) Training high quality annotators is difficult and time consuming. (2) Using solely human annotation for monitoring is impractical to support millions of daily posts. To this end, it is desirable to have intelligent systems which can perform well at the defined classification task across a plethora of FQ categories, preferably requiring little to no training. Hence with the recent releases of LLMs equipped with the ability to process multimodal inputs, it is natural to question whether we can utilize these off the shelf models for our task. Particularly, we are interested in the performance of GPT-4o at classifying TikTok videos according to the defined text policies, treating it as black-box classifier, with no additional training. Such a result would not only provide us with insight on GPT-4o’s vision-language understanding, but also serve as a comparison for application specific models currently employed, as well as a baseline for more costly approaches requiring additional training.
3.2 Method
In this section, we detail how we implement GPT-4o as a vision-language classifier for FQ categories in our experiments, and refer the reader to the official GPT-4o API used for reference OpenAI (2025). We provide a high level overview of our experiment design in Figure 1. Given a FQ category defined by the text strings {CATEGORY}, the name of the category, and {POLICY}, the detailed text policy for the corresponding category, we provide GPT-4owith the following system level instructions:
## Task: Label Videos for {CATEGORY}Content ### Objective: You are required to classify videos based on if they are {CATEGORY}content. The classification will help in training models to identify {CATEGORY}content. Each video should be labeled with an associated score indicating the likelihood of the video being {CATEGORY}content. ### Detailed Policy: #### {CATEGORY}: {POLICY} ### Output: For each video, output a clear reasoning behind your decision and a score (0-100) indicating the overall likelihood of the video being {CATEGORY}content. Format your output as a JSON object with the following keys: reasoning: a chain of reasoning that explains how you arrived at your classification. score: An integer score from 0 to 100 representing how likely it is that the video is {CATEGORY}content, with 100 indicating that the video certainly is {CATEGORY}content, and 0 indicating that you are confident this is not {CATEGORY}. ### Notes: - Be clear and specific in your classification. - Inspect every frame provided. - Use the detailed policy to guide your judgment.
@@ 3.2 Method
Given a video with the following features: audio transcription: {asr} hashtag: {hashtag} text: {text} sticker text: {stickerText} video frames: images with base64 encoding provided Format your output as a JSON object with the specified keys.
@@ 4.1 Tasks
We conducted experiments using OpenAI’s latest GPT-4o model, accessed through Microsoft Azure at the time of the study. The temperature parameter was set to zero, with the top $p=1$ value fixed at one, without employing any stop-words. Additionally, the frequency penalty was set to 0.5, while the presence penalty remained at zero. The Gemini version we compare in our preliminary study is Gemini-1.5-pro. For each video, we sampled frames at $0.5$ frames per second, including the first and last frames, and used a maximum of 30 frames irrespective of the video length.
@@ 4.2 Baselines
For the domain specific tasks we called out earlier, we build classification models that leverages multi-modal features and human annotation results for training as the baselines. The major difference between the baseline models and GPT-4o is baselines models are either rule-based or trained with annotated samples, while the GPT-4o is pretrained on large vision-language corpora with a larger model size. Specifically, for SF/Non-Interactive/Watermark, we could not collect enough training data to train a classifier by the time of this experiment. We tried rule-based method on these issues and only got acceptable performance on SF. We omit reporting the baseline performance for Non-Interactive/Watermark in our experiment.
@@ 4.4 Analysis
We analyze our research by answer the following $6$ questions sequentially, utilizing the aforementioned experiments.
@@ 5 Conclusion
In this paper, we investigate the application of GPT-4o on video classification and provide a strong baseline on the application of LLMs in video classification and other multimodal tasks. By demonstrating the potential of generative models in solving practical industry problems, we offer insights that can guide the development of more effective, scalable solutions for industries that rely on multimodal data. Moreover, our findings challenge the notion that LLMs are not suited for industry-specific applications, showing that with the right adaptations, these models can be both powerful and practical tools for real-world classification problems.
Limitations: We do not consider finetuning, as our focus is one shot performance. Additionally, we do not list results of other VLM models (e.g. Gemini), but found that initial experiments showed very similar performance between them.

### v3

@@ Abstract
In this study, we tackle industry challenges in video content classification by exploring and optimizing GPT-based models for zero-shot classification across seven critical categories of video quality. We contribute a novel approach to improving GPT’s performance through prompt optimization and policy refinement, demonstrating that simplifying complex policies significantly reduces false negatives. Additionally, we introduce a new decomposition-aggregation-based prompt engineering technique, which outperforms traditional single-prompt methods. These experiments, conducted on real industry problems, show that thoughtful prompt design can substantially enhance GPT’s performance without additional finetuning, offering an effective and scalable solution for improving video classification.
@@ 1 Introduction
In the past decade, recognition over mutliple modalities has become an increasingly critical challenge for large video platform companies, such as TikTok and YouTube, which rely on managing vast amounts of user-generated content. Unlike traditional classification tasks that process single-modality inputs, multi-modal classification combines multiple data sources, such as image, audio, and text, to provide more accurate and context-aware predictions Wang et al. (2017). The ability to reason across these different modalities is essential for effective content classification. This capability is particularly important for identifying inappropriate content, or videos that do not align with platform policies, as it ensures a more nuanced understanding of content beyond just visual or textual cue de Freitas et al. (2019); Yousaf & Nawaz (2022).
@@
With the advent of large language models (LLMs) like GPT-4, a new paradigm has emerged in the field of multi-modal classification Achiam et al. (2023). LLMs offer the promise of generalization, boasting the ability to perform a wide range of tasks with minimal task-specific data, often through one-shot or few-shot learning. Their large-scale pretraining on diverse datasets imbues them with a form of “world knowledge” that can be flexibly adapted to a range of problems, including those that involve multi-modal inputs. As a result, generative AI models are now being explored as a promising solution for problems traditionally addressed by multi-modal architectures Tang et al. (2023). One domain ripe for this exploration is video classification, where the challenge lies in understanding video content through a combination of visual, auditory, and sometimes textual cues.
@@
However, while LLMs like GPT-4 offer broad generalization, applying them to real-world, industry-specific problems has proven difficult. These models are typically pretrained on massive public datasets, meaning that company-specific data—especially multi-modal data related to proprietary video content—is rarely seen during pretraining. As such, LLMs may struggle to capture the nuances of these specific tasks without further tuning Prottasha et al. (2024); Zheng et al. (2024); Wang et al. (2024). Additionally, despite the hype surrounding generative models, there is still skepticism in the industry regarding their return on investment Sachs (2024).
@@
In this study, we aim to address these challenges by evaluating the performance of GPT-4 in real-world, industry-relevant video classification tasks that require multi-modal understanding. Specifically, we explore the potential of GPT-4 for classifying TikTok video content based on a variety of criteria. Our work differs from traditional multi-modal models in two key ways: first, by focusing on the zero-shot and few-shot capabilities of GPT-4, we investigate whether large, pretrained generative models can be effectively applied to multi-modal tasks without extensive fine-tuning. Second, we explore the impact of policy refinement and prompt engineering on GPT-4’s performance, addressing one of the core limitations of LLMs in industry—namely, their difficulty in handling complex, highly specific policies like those used to moderate video content.
@@
Our contributions are threefold. First, we present a comprehensive evaluation of GPT-4’s ability to handle real-world, multi-modal video classification tasks, comparing its performance to existing, in-house classification models that leverage multi-modal features. We show that GPT-4 can perform on par with these models in several categories, but faces challenges in more complex classifications, such as clickbait videos. Second, we introduce novel techniques to enhance GPT-4’s performance in these tasks. We find that shortening complex policy prompts improves GPT-4’s ability to classify videos accurately by reducing false negatives, and that prompt engineering—specifically, dividing tasks like clickbait detection into subcategories—can lead to significant performance gains. Lastly, our experiments are grounded in actual industry datasets and problems, making the results highly applicable to video classification tasks at scale.
@@ 2 Related Work
In this section, we review topics related to our study, including general and generative multi-modal models.
Multi-modal Classification. Historically, multi-modal modeling has been approached using modality-specific architectures. For instance, BERT (Devlin, 2018), RoBERTa (Liu, 2019), ResNet (He et al., 2016), and ViT (Dosovitskiy, 2020) are commonly employed for text and vision tasks, respectively. Multi-modal models like CLIP (Radford et al., 2021) and GLIP (Li et al., 2022) have effectively combined visual and textual inputs, leveraging correlations between modalities to achieve strong classification and detection results. However, these models typically require extensive task-specific fine-tuning and are limited by the rigid nature of their underlying architectures. This has resulted in a fragmented landscape where models are highly specialized for narrow tasks and lack generalizability to unseen problems or broader domains without significant retraining.
@@ 3.1 Problem Setup
Two major challenges to this problem are: (1) Training high quality annotators is difficult and time consuming. (2) Using solely human annotation for monitoring is impractical to support millions of daily posts. To this end, it is desirable to have intelligent systems which can perform well at the defined classification task across a plethora of FQ categories, preferably requiring little to no training. Hence with the recent releases of LLMs equipped with the ability to process multi-modal inputs, it is natural to question whether we can utilize these off the shelf models for our task. Particularly, we are interested in the performance of GPT-4 at classifying TikTok videos according to the defined text policies, treating it as black-box classifier, with no additional training. Such a result would not only provide us with insight on GPT-4o’s vision-language understanding, but also serve as a comparison for application specific models currently employed, as well as a baseline for more costly approaches requiring additional training.
3.2 Formulation
In this section, we detail how we implement GPT-4 as a vision-language classifier for FQ categories in our experiments, and refer the reader to the official API used for reference OpenAI (2025). We provide a high level overview of our experiment design in Figure 1. Given a FQ category defined by the text strings {CATEGORY}, the name of the category, and {POLICY}, the detailed text policy for the corresponding category, we provide GPT-4o with the following system level instructions:
  ## Task: Label Videos for {CATEGORY}Content
### Objective:
You are required to classify videos based on if they are {CATEGORY}content. The classification will help in training models to identify {CATEGORY}content. Each video should be labeled with an associated score indicating the likelihood of the video being {CATEGORY}content.
### Detailed Policy:
#### {CATEGORY}:
{POLICY}
### Output:
For each video, output a clear reasoning behind your decision and a score (0-100) indicating the overall likelihood of the video being {CATEGORY}content.
Format your output as a JSON object with the following keys:
reasoning: a chain of reasoning that explains how you arrived at your classification.
score: An integer score from 0 to 100 representing how likely it is that the video is {CATEGORY}content, with 100 indicating that the video certainly is {CATEGORY}content, and 0 indicating that you are confident this is not {CATEGORY}.
### Notes:
- Be clear and specific in your classification.
- Inspect every frame provided.
- Use the detailed policy to guide your judgment.
@@ 3.2 Formulation
  Given a video with the following features:
     audio transcription: {asr}
     hashtag: {hashtag}
     text: {text}
     sticker text: {stickerText}
     video frames: images with base64 encoding provided
Format your output as a JSON object with the specified keys.
@@ 4 Results
Before providing a detailed analysis of the results, we give additional details on the setup of our experiments.
@@ 4.1 Tasks
We conducted experiments using OpenAI’s latest GPT-4o model, accessed through Microsoft Azure at the time of the study: gpt-4o version 2024-05-13. While we also used Gemini-1.5-pro during preliminary experiments, we found the performance was similar and left it out when computing our final results. The temperature parameter was set to zero, with the top $p=1$ value fixed at one, without employing any stop-words. Additionally, the frequency penalty was set to 0.5, while the presence penalty remained at zero. For each video, we sampled frames at $0.5$ frames per second, including the first and last frames, and used a maximum of 30 frames irrespective of the video length.
@@ 4.2 Baselines
For the aforementioned domain specific tasks, we used classification models trained on multi-modal features and human annotations as the baselines. The major difference between these baseline models and GPT-4o is that the baseline models are either rule-based or trained on the annotated samples, while GPT-4o is a much bigger model that is pretrained on large vision-language corpora. Additionally, for SF/Non-Interactive/Watermark, we could not collect enough training data to train a classifier by the time of this experiment. We tried rule-based method on these issues and only got acceptable performance on SF. Therefore, we omit reporting the baseline performance for Non-Interactive/Watermark in our experiment.
@@ 4.4 Analysis
We analyze our research by answering the following $6$ questions sequentially, utilizing the aforementioned experiments.
@@ 5 Conclusion
In this paper, we investigate the application of GPT-4 on video content moderation, providing a strong baseline for the use of LLMs in multi-modal classification. By demonstrating the potential of generative models in solving this industry problem, we offer insights that can guide the development of more effective, scalable solutions for industries that rely on classifying multi-modal data. Our findings show that with the right adaptations, these models can be powerful and practical tools for industry-specific classification problems.
Limitations: We could not test against new reasoning-models such as Deepseek’s R1 or OpenAI’s o1 and o3 models as they were not released at the time of this study. However, we believe that their increased capabilities do not outweigh the larger cost when considering the scale of content moderation in industry. Furthermore, prompt-engineering techniques, like the one introduced in Exp. 3, can be seen as a cost-effective way to inject reasoning that is calibrated for the task at hand.
Future Directions: A clear direction for future works is finetuning. While performance increase is one benefit, inference cost is also important for industry applications. Therefore, clever design that integrates vision and text modalities with smaller, fine-tuned LLM’s, is a promising direction. Furthermore, audio should be considered as an additional modality once pre-trained foundation models are capable of processing it alongside video and text.